# VTAO-BiManip: Masked Visual-Tactile-Action Pre-training with Object Understanding for Bimanual Dexterous Manipulation

### v1

@@ Abstract
Bimanual dexterous manipulation remains significant challenges in robotics due to the high DoFs of each hand and their coordination. Existing single-hand manipulation techniques often leverage human demonstrations to guide RL methods but fail to generalize to complex bimanual tasks involving multiple sub-skills. In this paper, we introduce VTAO-BiManip, a novel framework that combines visual-tactile-action pretraining with object understanding to facilitate curriculum RL to enable human-like bimanual manipulation. We improve prior learning by incorporating hand motion data, providing more effective guidance for dual-hand coordination than binary tactile feedback. Our pretraining model predicts future actions as well as object pose and size using masked multimodal inputs, facilitating cross-modal regularization. To address the multi-skill learning challenge, we introduce a two-stage curriculum RL approach to stabilize training. We evaluate our method on a bottle-cap unscrewing task, demonstrating its effectiveness in both simulated and real-world environments. Our approach achieves a success rate that surpasses existing visual-tactile pretraining methods by over 20%.
Figure 1: Left: examples of our bottle-cap turning dataset. Middle: simulation results of the bottle-cap task with Shadow Hand [1]. Right: the real-world deployment of the bottle-cap task with Leap Hand [2].
1College of Control Science and Engineering, Zhejiang University, Hangzhou, 310027, ChinaQi Ye (Corresponding author, qi.ye@zju.edu.cn) is with the College of Control Science and Engineering and the State Key Laboratory of Industrial Control Technology, Zhejiang University, and also with the Key Key Lab of CS&AUS of Zhejiang Province.This work was supported in part by the National Natural Science Foundation of China (Grant Number: 62088101, 62103372, 62233013).
@@ I Introduction
Humans can perform various complex bimanual tasks in daily life while enabling robots to learn human-like bimanual manipulation skills is challenging due to the high degrees of freedom (DoFs) of each hand, and the coordination between hands. An algorithm needs to search in the high dimensional space for a two-hand configuration that conforms to physical laws and also meets the task requirements.
Solving a similar problem, existing work on single high DoFs robotic hand manipulation [3, 4, 5, 6] leverage priors from human demonstrations to tackle the challenge. In M2VTP [6], the prior knowledge of the human manipulation skill of bottle-cap turning is learned by a visual-tactile pre-training network and incorporated into a reinforcement learning (RL) method. However, the task in [6] involves turning a bottle cap with the bottle fixed on a table. In our task, the bottle is required to be grasped by the left hand and held in the air while the right hand performs bottle-cap turning as shown in Fig. 1 (middle). Considering the task similarity, the visual-tactile pre-training and RL method [6] is then applied to our task by collecting the human bimanual manipulation data and changing the state and action according to two hands.
However, the method failed in the RL training stage. The curves for the rewards and success rate stay near zero after thousands of iterations. The reasons can be attributed to two factors. Our task consists of three sub-tasks (grasp, hold, and bottle-cap turning). Learning the three skills and the coordination between these skills and two hands is required. 1) In M2VTP [6], the extra binary tactile modality helps the coordination between the fingers of a single hand in the bottle-cap turning by pretraining, but it cannot help the coordination between these skills and two hands as that much since touch status does not change in many segments. 2) When applying the RL method in [6] to our task, learning all these skills and coordination is treated as a black box and it is hard to learn all of them together.
To address the issues, in this paper, we propose VTAO-BiManip, a novel pretraining and RL learning framework for bimanual manipulation. It learns priors from more modalities during human manipulation and adopts a curriculum reinforcement learning strategy. Specifically, for the first issue, we observe that 1) Unlike binary tactile information, which only indicates physical contact, human hand actions offer richer and more intuitive priors on movement and coordination, effectively guiding downstream bimanual tasks; 2) the size and pose of objects determine how the hands must adjust and coordinate during manipulation, incorporating their perception is essential for achieving accurate and adaptable bimanual tasks. To learn these two priors, we add the action and object pose ’modality’ upon the visual-tactile pretraining network as shown in  Fig. 3. Notice that different from imitation learning which predicts the next action based on the input state, our action pretraining predicts successive future actions with masked input hand state. The masked and future action prediction strategy allows the pretraining network to benefit from a form of ’cross-modal regularization’ in learning the action prior due to the presence of other modality data. To pretrain our approach VTAO-BiManip, we use the same collection system of [6] and add a motion capture system to collect human visual-tactile manipulation data, hand movements, and object pose information.
@@
For the second issue, a two-stage curriculum reinforcement learning framework is proposed: in the first stage, the bottles are fixed on the table, encouraging the left hand to grab the bottle and the right hand to unscrew the cap; in the second stage, the bottle is released, forcing the two-hand cooperative manipulation skill learning.
@@
  • We propose VTAO-BiManip, a novel visual-tactile-action pretraining method with object understanding for bimanual dexterous manipulation. This approach leverages human data to pretrain models for action prediction and object state estimation.
@@
  • We pretrain VTAO-BiManip with collected human demonstrations and deploy to the downstream bimanual cap unscrewing task both in simulation and the real world.
@@ II-A Multimodal Pretraining for Robotics
Pre-training methods utilizing unsupervised learning [7, 8, 9] have emerged as particularly effective in multimodal representation generalization. Due to the availability of extensive manipulation datasets that include visual and language modalities [10, 11, 12], some researches [13, 14, 15, 16] explores the representation pretraining of these modalities in robotic tasks, demonstrating their effectiveness and generalization capabilities.  [17] extends this by integrating action modalities with vision and language, enabling robots to predict and execute tasks. [18] improves downstream performance by predicting future motions of any point via visual pertaining. However, pretraining on action modalities has primarily focused on simple end-effectors, with limited exploration in dexterous manipulation. For tactile, studies such as [19, 20, 21, 6] highlight the effectiveness of visuo-tactile fusion in robotic manipulation. Yet, its application in bimanual tasks remains underexplored. To address this gap, we propose a novel approach, incorporating multimodal pretraining, including action prediction, into bimanual manipulation.
@@ II-B Robotic Bimanual Manipulation
Learning-based bimanual manipulation has gained significant attention in robotics. While some studies have successfully utilized demonstration data [22, 23, 24], the lack of high-quality multi-fingered hand demonstrations often restricts their applicability to simple end-effectors. Efforts to collect such data for multi-fingered hands have been made [25, 26, 27], but challenges such as retargeting errors and latency limit their practical use.
On the other hand, sim-to-real approaches have also attracted growing interest in the domain of robotic manipulation [28, 29, 30, 31], demonstrating numerous successes across various tasks such as deformable object manipulation [32], open doors [33], and in-hand rotations [34]. However, most of these methods focus on either single dexterous hands [35, 36, 37, 38] or dual-gripper systems [39, 40]. Some progress has been made in dual-dexterous-hand manipulation [41, 42, 43].  [41] construct a dexterous bimanual manipulation platform, offering 20 manipulation tasks.  [42] employ multi-agent RL to train a bimanual handover task, incorporating a trajectory prediction model to bridge the sim-to-real gap.  [43] presents a system where robot hands are trained to play the piano using DRL. However, these works primarily rely on visual and proprioceptive modalities, which somewhat limit the performance of manipulation tasks. Multimodal data has proven effective in dual-gripper [44, 20, 45] and single dexterous hand tasks [46, 47, 48], its application in bimanual dexterous manipulation remains largely unexplored. Our work represents a pioneering effort in applying multimodal sensing to dual-dexterous-hand manipulation.
@@ III Method
In this section, we first introduce our data collection system for capturing ego-centric human manipulation data, including Vision, Tactile, Action, and Object information (VTAO), detailed in Section III-A. Section III-B presents VTAO-BiManip, a novel pretraining framework that utilizes human manipulation priors for multimodal feature extraction and fusion. Finally, Section III-C outlines the development of a bimanual dexterous bottle-cap unscrewing task in simulation and the subsequent policy training using RL.
@@ III-A VTAO Capture System for Human Bi-Manipulation
We develop a novel VTAO data collection system for capturing multimodal data during human bimanual manipulation, built upon our previous visual-tactile acquisition system [6].
@@ III-A1 System Overview
The VTAO capture system includes: 1) tactile and motion acquisition gloves for both hands, 2) Hololens2  [^1] https://www.microsoft.com/en-us/hololens/buy for visual data capture, 3) the ChingMu motion capture (mocap) system  [^2] https://www.chingmu.com/ for recording hand movements and 6-DoF object pose, and 4) a personal computer responsible for data acquisition and alignment, as shown in Fig. 2. Each acquisition glove consists of two layers: an inner layer with 20 tactile sensors, as [6], and an outer layer with markers for motion capture. The mocap system comprises 16 cameras to capture the 6-DoF pose information of the bottle and the precise movements of the hand joints.
@@ III-A2 Data Alignment and Unified Representation
Multimodal data are captured independently, with visual data sampled at 30 Hz, tactile data at 200 Hz, and mocap data at 1 kHz. Data from other modalities are synchronized with the visual modality using local timestamps, which are recorded by the personal computer during the acquisition. To unify coordinate systems across modalities, we place markers on the Hololens2 and calibrate the transformation between the camera and the rigid body coordinates of the Hololens2. Utilizing this transformation allows us to map the 6-DoF pose of the bottle, as well as the hand motion data, to the RGB camera coordinate system of the visual modality, facilitating model comprehension.
@@ III-A3 Retarget Human Manipulation Trajectories
To unify bimanual action representations across pretraining and downstream tasks, we retarget human bimanual actions from the dataset onto the dual dexterous hands used in downstream tasks. Following DexRepNet [29], we regard retargeting as a nonlinear optimization problem. Our objective is to ensure similarity in distances from fingertips and finger segments to the wrist, in both robotic and human hands at all stages, resulting in natural robotic hand postures. Our objective function aligns with that presented in [25], formulated as:
@@
where $\mathbf{v}_{i}^{\mathbf{R}}$ and $\mathbf{v}_{i}^{\mathbf{H}}$ signify the target vectors of the robotic and human hands respectively (from wrist to fingertips and wrist to finger segments). These are determined using forward kinematics of the dual Shadow Hand $M^{\mathbf{R}}$ and human hands $M^{\mathbf{H}}$, coupled with joint angles $q^{\mathbf{R}}_{t}$ and $q^{\mathbf{H}}_{t}$ at step $t$. $q_{t}^{\mathbf{R}*}$ denotes the optimized joint angles of the robotic hands at step $t$. We refine each step of the human bimanual manipulation trajectories within the dataset to generate retargeted robotic hand trajectories, which are subsequently utilized for the pre-training of the robots.
@@
Figure 3: Overview of our VTAO-BiManip pre-train framework and its downstream application.
III-B Masked VTAO Transformer for Pretraining
To explore the benefits of incorporating multimodal human manipulation data into dual-hand robotic skill learning, we propose a novel pre-training framework, VTAO-BiManip, based on MAE [8]. This framework uses our collected VTAO data for pre-training. It extracts and integrates latent environmental representations to support the learning of downstream robotic manipulation tasks. VTAO-BiManip includes a VTAO encoder and a VTAO decoder, as shown in Fig. 3 (left).
III-B1 VTAO Encoder $E_{\theta}(V,C,A)\rightarrow(h_{\text{cls}},h_{v},h_{c},h_{a},\allowbreak h_{o})$
The VTA encoder takes in the visual, tactile, and action pair (V, C, A) as input. The RGB image $V\in\mathbb{R}^{224\times 224\times 3}$, the binarized tactile signal from both hands $C\in\mathbb{R}^{40}$ (thresholded by $\lambda$), and the joint angles $A\in\mathbb{R}^{48}$, retargeted to the dual Shadow Hand, are processed via a Conv2D and two MLPs to form a patch sequence. Additionally, empty patches are added to align feature dimensions. These patches are used by the decoder to reconstruct the bottle information $\widehat{O}$. After adding positional encoding, the patch sequence becomes a token sequence. Following MAE [8], we randomly mask $r_{v}\in[0,1)$ of the visual tokens, $r_{c}\in[0,1)$ of the tactile tokens, $r_{a}\in[0,1)$ of the action tokens, and $r_{o}\in[0,1)$ of the empty tokens to form the input token sequence $(t_{v},t_{c},t_{a},t_{\text{null}})$. A CLS token $t_{\text{cls}}$ is inserted into this sequence as an aggregated representation for downstream tasks. The final token sequence $(t_{\text{cls}},t_{v},t_{c},t_{a},t_{\text{null}})$ is processed by a Transformer [24] encoder, producing a latent representation sequence $(h_{\text{cls}},h_{v},h_{c},h_{a},h_{o})$.
III-B2 VTAO Decoder $D_{\theta}(h_{v},h_{c},h_{a},h_{o},m)\rightarrow(\widehat{V},\widehat{C},\allowbreak[\widehat{A},\widehat{A_{+1}},...,\widehat{A_{+p}}],\widehat{O})$

### v2

@@ Abstract
Bimanual dexterous manipulation remains a significant challenge in robotics due to the high DoFs of each hand and their coordination. Existing single-hand manipulation techniques often leverage human demonstrations to guide RL methods but fail to generalize to complex bimanual tasks involving multiple sub-skills. In this paper, we propose VTAO-BiManip, a novel framework that integrates visual-tactile-action pre-training with object understanding, aiming to enable human-like bimanual manipulation via curriculum reinforcement learning (RL). We improve prior learning by incorporating hand motion data, providing more effective guidance for dual-hand coordination. Our pretraining model predicts future actions as well as object pose and size using masked multimodal inputs, facilitating cross-modal regularization. To address the multi-skill learning challenge, we introduce a two-stage curriculum RL approach to stabilize training. We evaluate our method on a bimanual bottle-cap twisting task, demonstrating its effectiveness in both simulated and real-world environments. Our approach achieves a success rate that surpasses existing visual-tactile pretraining methods by over 20%.
Figure 1: Left: examples of our bottle-cap turning dataset. Middle: simulation results of the bottle-cap task with Shadow Hand [shadowhand]. Right: the real-world deployment of the bottle-cap task with Leap Hand [shaw2023leaphand].
1College of Control Science and Engineering, Zhejiang University, Hangzhou, 310027, ChinaQi Ye (Corresponding author, qi.ye@zju.edu.cn) is with the College of Control Science and Engineering and the State Key Laboratory of Industrial Control Technology, Zhejiang University, and also with the Key Key Lab of CS&AUS of Zhejiang Province.This work was supported in part by NSFC under Grants (No. 62088101, 62233013, 62293511) and Key Research and Development Program of Zhejiang Province (No. 2025C01072).
@@ I Introduction
Humans effortlessly perform diverse complex bimanual tasks like opening containers and assembling furniture in daily life via dynamic inter-hand coordination. Such capabilities rely on three fundamental components: 1) kinematic inter-hand coordination: dynamically adjusting joint-level movements between hands to maintain operational stability, 2) real-time object-state sensing: modifying manipulation strategies based on real-time state changes of manipulated objects, 3) temporally composition of complementary skill primitives: executing phase-dependent primitives (grasping, holding, and rotating) under physical constraints. While humans perform these tasks effortlessly, robotic dexterous bimanual manipulation remains an open challenge. The control complexity grows exponentially when coordinating two high-DoF hands (totaling 30+ joints) under physical constraints, particularly in dynamic scenarios with coupled dynamics between object states and manipulation requirements.
Recent advances in dexterous manipulation [Rajeswaran2018DAPG, qin2022dexmv, wu2022ILAD, liu2024m2vtp] have demonstrated progress in single-hand scenarios by leveraging human demonstrations and multi-modal sensing. Particularly, Liu et al. [liu2024m2vtp] established a visual-tactile pre-training framework that enables transferable manipulation skills through masked autoencoding. However, these methods fundamentally rely on constrained task setups - for instance, fixing objects on tables to bypass bimanual coordination challenges. Directly extending such frameworks to dynamic bimanual tasks like bottle-cap twisting reveals three irreducible challenges: 1) kinematic coordination limitation: while vision and binary tactile sensing aids single-hand finger coordination [liu2024m2vtp], they fail to capture inter-hand kinematic synergy critical for dynamic bimanual manipulation. 2) object-state decoupling: when the bottle tilts during cap rotation, the passive hand must respond based on real-time 3D pose changes. Current methods lack explicit perception of object pose and size, leading to incoherent hand movements under dynamic states. 3) sub-skill composability: successful cap twisting requires temporally aligned sub-skills - the passive hand must establish stable grasping before the active hand initiates rotation. However, single-stage policies in prior work [liu2024m2vtp] lack mechanisms to coordinate these interdependent sub-processes.
To address the issues, in this paper, we propose VTAO-BiManip, a novel framework extending [liu2024m2vtp] through two key innovations: visual-tactile-action pre-training method with object understanding for multi-modal perception and curriculum reinforcement learning for action control policy learning. To address kinematic coordination blindness and object-state decoupling, we introduce hand action and object-centric understanding as complementary modalities during pre-training: 1) while prior work relies on visual-tactile fusion, we introduce hand joint angles as explicit kinematic observables and predict future actions to model inter-hand synergy. This enables the policy to dynamically coordinate bimanual motions; for instance, progressively adjusting the finger joint angles of the passive hand to counteract active-hand torque during cap twisting. Notably, unlike imitation learning which predicts next-step actions from current states, our action pre-training predicts future action sequences using masked multi-modal inputs. 2) we also augment the pre-training framework with object pose and size estimation, forcing the encoder to learn geometry-aware representations. By reconstructing the bottle’s 6D pose and 3D dimensions from visual-tactile-action inputs, the model gains explicit awareness of dynamic state changes. During deployment, this allows real-time adaptation.
@@
To address sub-skill composability, a two-stage curriculum reinforcement learning framework is proposed. In the first stage, the bottles are fixed on the table, encouraging the left hand to grab the bottle and the right hand to unscrew the cap. This phase reduces exploration complexity, building foundational sub-skills of both hands. In the second stage, the bottle is physically released, forcing the two-hand cooperative manipulation skill learning.
@@
  • We propose VTAO-BiManip, a novel visual-tactile-action pre-training method with object understanding for bimanual dexterous manipulation. This approach leverages human data to pre-train models for action prediction and object state estimation.
@@
  • We propose a bimanual bottle-cap twisting task where our VTAO-BiManip pre-training and curriculum RL framework succeeds both in simulation and the real world.
@@ II-A Multimodal Pretraining for Robotics
Pre-training methods that use unsupervised learning [radford2021clip, he2022mae, oord2018representation] have emerged as particularly effective in multimodal representation generalization. Due to the availability of extensive manipulation datasets that include visual and language modalities [goyal2017sth_sth_v2, grauman2022ego4d, Damen2022epic_kitchen], some researches [ma2023liv, nair2022r3m, karamcheti2023voltron, xiao2022mvp] explore the pretraining representation of these modalities in robotic tasks, demonstrating their effectiveness and generalization capabilities.  [brohan2023rt2] extends this by integrating action modalities with vision and language, enabling robots to predict and execute tasks. [wen2023atm] improves downstream performance by predicting future motions of any point via visual pertaining. However, pre-training on action modalities has primarily focused on simple end-effectors, with limited exploration in dexterous manipulation. For tactile, studies such as [chen2023VTT, mejia2024hearing, gano2024vt, liu2024m2vtp] highlight the effectiveness of visuo-tactile fusion in robotic manipulation. However, its application in bimanual tasks remains underexplored. To address this gap, we propose a novel approach that incorporates multimodal pre-training, including action prediction, into bimanual manipulation.
@@ II-B Robotic Bimanual Manipulation
Learning-based bimanual manipulation has gained significant attention in robotics. Although some studies have successfully used demonstration data [stepputtis2022system, zhao2023aloha, kim2021transformer], the lack of high-quality multi-fingered hand demonstrations often restricts their applicability to simple end-effectors. Efforts to collect such data for multi-fingered hands have been made [handa2020dexpilot, qin2022one, arunachalam2023dexterous], but challenges such as retargeting errors and motion retargeting latency limit their practical use.
Sim-to-real approaches have demonstrated success in robotic manipulation [akkaya2019solving, liu2023dexrepnet, qin2023dexpoint, qi2023hand], including deformable object manipulation [matas2018sim], open doors [rajeswaran2017DAPG], and in-hand rotations [chen2023visual_dexterity]. However, most methods focus on single dexterous hands [chen2023sequential, yin2023RotateWithoutSeeing, bao2023dexart, pitz2023dextrous] or dual-gripper systems [li2023efficient, lin2023bi_touch]. Some progress has been made in dual-dexterous-hand manipulation [chen2023bi, huang2023dynamic_handover, zakka2023robopianist, Lin2024twisting][chen2023bi] construct a dexterous bimanual manipulation platform, offering 20 manipulation tasks. [huang2023dynamic_handover] employ multi-agent RL to train a bimanual handover task, incorporating a trajectory prediction model to bridge the sim-to-real gap. [zakka2023robopianist] presents a system in which robot hands are trained to play the piano using DRL. [Lin2024twisting] explores RL-based twisting policies for predefined cylindrical containers using sparse visual keypoints, demonstrating limited zero-shot transfer through contact-aware rewards. However, these works mainly rely on visual and proprioceptive modalities, which somewhat limit the performance of manipulation tasks. Multimodal data has proven effective in dual-gripper [hogan2020tactile, mejia2024hearing, sunil2023visuotactile_cloth] and single dexterous hand tasks [yuan2024RobotSynesthesia, qi2023inhand-visuotactile, guzey2024see], its application in bimanual dexterous manipulation remains largely unexplored. Our work represents a pioneering effort to apply multimodal sensing to dual-dexterous-hand manipulation.
@@ III Method
In this section, we first introduce our data collection system to capture ego-centric human manipulation data, which includes Vision, Tactile, Action, and Object (VTAO) information (see Section III-A). Section III-B then introduces VTAO-BiManip, a novel pre-training framework that utilizes human manipulation priors for multimodal feature extraction and fusion. Finally, Section III-C details the implementation of a dexterous bimanual bottle cap unscrewing task in simulation, followed by policy training through the curriculum RL.
@@ III-A VTAO Capture System for Human Bi-Manipulation
We develop a VTAO data collection system to capture multimodal data during human bimanual manipulation. The visual and tactile acquisition methods are directly adopted from our previous system [liu2024m2vtp]. The key addition is a motion capture system that records hand motion trajectories and bottle cap poses in the same coordinate system. Below, we detail this VTAO capture system:
@@ III-A1 System Overview
The VTAO capture system includes: 1) Hololens2  [^1] https://www.microsoft.com/en-us/hololens/buy for visual data capture, 2) tactile and motion acquisition gloves for both hands, 3) the CHINGMU motion capture system  [^2] https://www.chingmu.com/ with 16 cameras to capture precise hand movements and 6-DoF object pose, and 4) a personal computer responsible for data acquisition and alignment, as shown in Fig. 2. The CHINGMU system plays a critical role in acquiring high-quality bimanual motion trajectories and object pose data, which are essential for the following model pre-train. Each acquisition glove consists of two layers: an inner layer with 20 tactile sensors (identical to the design in [liu2024m2vtp]), and an additional outer layer with retroreflective markers for optical motion capture.
@@ III-A2 Data Alignment and Unified Representation
Multimodal data are captured independently, with visual data sampled at 30 Hz, tactile data at 200 Hz, and mocap data at 1 kHz. Data from other modalities are synchronized to the visual modality through local timestamps recorded by the personal computer during acquisition. To unify the motion capture and visual coordinate systems, we install markers on both the Hololens2 and a calibration board, then simultaneously capture their positions in the motion capture system while acquiring the calibration board’s location in the visual system, thereby calibrating the transformation between the camera coordinates and the Hololens2’s rigid-body coordinates. This transformation enables mapping of the 6-DoF bottle pose and hand motion trajectories to the RGB camera coordinate system of the visual modality, to facilitate model comprehension.
@@ III-A3 Retarget Human Manipulation Trajectories
To unify bimanual action representations across pre-training and downstream tasks, we retarget human bimanual trajectories to dexterous robotic hands via nonlinear optimization, following DexRepNet [liu2023dexrepnet]. The objective, aligned with [handa2020dexpilot], preserves geometric similarity by minimizing distances between wrist-to-fingertips and wrist-to-phalanxes across all motion phases, ensuring natural robotic hand poses, and is defined as:
@@
where $\mathbf{v}_{i}^{\mathbf{R}}$ and $\mathbf{v}_{i}^{\mathbf{H}}$ denote the target vectors for the robotic and human hands respectively, defined as wrist-to-fingertip and wrist-to-phalanx distance vectors. These vectors are computed via forward kinematics of the dual Shadow Hand $M^{\mathbf{R}}$ and human hand model $M^{\mathbf{H}}$, using joint angles $q^{\mathbf{R}}_{t}$ and $q^{\mathbf{H}}_{t}$ at timestep $t$. The optimized joint angles $q_{t}^{\mathbf{R}*}$ are obtained through this minimization process. We optimize each timestep of the collected human bimanual trajectories to generate retargeted robotic motion sequences, which are subsequently used for robotic policy pre-training.
@@
Figure 3: Overview of our VTAO-BiManip pre-train framework and its downstream bimanual manipulation with sub-skill and collaborative curriculum reinforcement learning.
III-A4 VTAO Dataset
Using our VTAO acquisition system, we collect 216 human bimanual manipulation trajectories with 26 different bottles, spanning 6-17 seconds per trajectory and totaling 61,684 temporally aligned multimodal frames. Through preprocessing, RGB images are cropped to $224\times 224$ resolution, tactile signals are binarized with a $\lambda=0.4V$ threshold, bottle poses are aligned to the camera coordinate system, and human hand motions are kinematically retargeted to dual robotic hands. Each processed frame includes a RGB image ($V\in\mathbb{R}^{224\times 224\times 3}$), bimanual tactile signals ($T\in\mathbb{R}^{40}$, 20 tactile sensors per hand), robotic joint angles of the current step ($A\in\mathbb{R}^{48}$, 24 joints per hand), and an object state vector ($O\in\mathbb{R}^{9}$) encoding the 6-DoF pose (position and orientation) and 3-D size information (width, height, depth) of the bottle.
III-B Masked VTAO Transformer for Pre-Training
To enhance dual-handed robotic skill learning through multimodal data fusion, we propose VTAO-BiManip, a MAE-based framework pre-trained on our collected VTAO dataset. As shown in Fig. 3, our pre-training framework follows an encoder-decoder structure: 1) a multimodal encoder processes visual, tactile, and current action inputs for each time step from the demonstration trajectories, projecting them into latent tokens that integrate information for different modalities via attention; 2) a decoder reconstructs the original sensory modalities, predicts subsequent actions, and estimates object states from these latent tokens. This joint reconstruction mechanism inherently enables cross-modal feature fusion during pre-training, where the model learns to recover complete environmental interactions from partial sensory observations. Notably, the object state information is only used for VTAO-BiManip pre-training and is not incorporated into RL policy learning during downstream tasks to maintain practical feasibility.
III-B1 VTAO Encoder $E_{\theta}(V,T,A)\rightarrow(h_{\text{fs}},h_{v},h_{c},h_{a},\allowbreak h_{o})$
The VTAO encoder processes a single frame input (V, T, A), comprising an RGB image $V$, binarized tactile signals $T$, and action $A$. The action $A$ is represented by robotic joint angles retargeted from human hand motions. Notice that the object information $O$ is excluded from inputs due to deployment constraints, but will be reconstructed via mask tokens appended to the input sequence. For all the input and output modalities, they are tokenized and each token is a vector of $d$ dimensions. The visual input ($V\in\mathbb{R}^{224\times 224\times 3}$) is divided into 196 non-overlapping patches via a Conv2D network to extract patch tokens $k_{v}\in\mathbb{R}^{196\times d}$. Binarized tactile signals $T\in\mathbb{R}^{40}$ (20 tactile sensors per hand) are projected into latent tokens $k_{c}\in\mathbb{R}^{40\times d}$ via an MLP, establishing a one-to-one mapping between tactile sensors and tokens. This binarization follows the design in [liu2024m2vtp], aiming to reduce the sim-to-real gap and eliminate the need for precise calibration across taxels. The action $A\in\mathbb{R}^{48}$ (joint angles from 24 joints per hand) is encoded into action tokens $k_{a}\in\mathbb{R}^{48\times d}$ through a separate MLP, with each token representing one joint angle. All input tokens are concatenated into a sequence $(k_{\text{fs}},k_{v},k_{c},k_{a},k_{\text{null}})$, where $k_{\text{fs}}\in\mathbb{R}^{d}$ is the fusion token (a special token aggregating global information for downstream tasks) and $k_{\text{null}}\in\mathbb{R}^{9\times d}$ are mask tokens for $O$ reconstruction. Following positional encoding and MAE-style masking [he2022mae], we randomly mask portions of tokens from different modalities: visual tokens $k_{v}$ (ratio $r_{v}\in[0,1)$), tactile tokens $k_{c}$ ($r_{c}\in[0,1)$), action joint tokens $k_{a}$ ($r_{a}\in[0,1)$) and object mask tokens $k_{\text{null}}$ ($r_{o}\in[0,1)$). The multi-modal tokens, with fusion token, are encoded by a Transformer [kim2021transformer], outputting latent representations $(h_{\text{fs}},h_{v},h_{c},h_{a},h_{o})$ for downstream tasks.
[…]